OpenStudy (anonymous):

Find the equation of the tangent line to the graph of the given function at the point with the indicated x-coordinate. f(x) = x + 2/x; x = 2

OpenStudy (anonymous):

\[f(x)=\frac{x+2}{x}\] \[f'(x)=\frac{-2}{x^2}\] \[f'(2)=-\frac{1}{2}\] is your slope, and your point is \[(2,2)\] and point slope formula give the answer

OpenStudy (anonymous):

Wait let me rewrite that equation.

OpenStudy (anonymous):

f(x) = x + (2/x); x = 2

OpenStudy (anonymous):

well... same idea, different answer

OpenStudy (anonymous):

My answer was y=-x+5 but it was incorrect.

OpenStudy (anonymous):

f prime of x = 1 + (-2/x^2) right?

OpenStudy (anonymous):

\[f'(x)=1-\frac{2}{x^2}\] \[f'(2)=1-\frac{1}{2}=\frac{1}{2}\] now use point slope with \[(2,3)\]

OpenStudy (anonymous):

Oh I see. I plugged in one instead of 2.

OpenStudy (anonymous):

.5x+4?